OpenStudy (anonymous):

The price of bronze has increased by 10% per year from 2000. In the year 2000, Harry bought a bronze medal for $120. Which of the following functions f(x) can be used to represent the price of the medal x years after 2000?

OpenStudy (anonymous):

(A) f(x) = 120(0.9)^x (B) f(x) = 10(120)^x (C) f(x) = 120(1.1)^x (D) f(x) = 10(1.1)^x

OpenStudy (anonymous):

Is this A?

OpenStudy (anonymous):

or C?

OpenStudy (anonymous):

@SolomonZelman @ganeshie8 @Godlovesme @hartnn Please I need Help!